"5 years ago mother was 7 times as old as her doughter. 5 years hence she will be 3 times as old as her doughter. find their present age?"
Answer by jorel1380(3719)

m-5=7(d-5)
m+5=3(d+5)
-------------
m-5=7d-35
m=7d-30
7d-30+5=3d+15
7d-25=3d+15
4d=40
d=10
m=40
The mother is 40; the daughter is 10..